OpenStudy (anonymous):

-5 - 2(3x - 4) = 21 ?

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

-5-6x+8=21 is the first step

OpenStudy (anonymous):

first youd have to do whats in the parenteses, mulitply 2 and 3, and 2 and 4 to get what he just said^

OpenStudy (anonymous):

already did that.

OpenStudy (anonymous):

then youd subtract -5 from both sides.

OpenStudy (anonymous):

just kidding youd add

OpenStudy (anonymous):

-6x+3=21 -6x=18 x=-3

OpenStudy (anonymous):

there you go..BEST OF LUCK ...YOU ROCK!!

OpenStudy (anonymous):

yeah dont listen to me im still a little rusty on multi step equations lol

OpenStudy (anonymous):

^naaah u r pretty good at this

OpenStudy (anonymous):

haha you're better than me @Ayeeitshannaahhh thanks! and thank you @dhirensharma

OpenStudy (anonymous):

anytime dont frget to fan so i can help you out

OpenStudy (anonymous):

haha i might need it right now lol again!

OpenStudy (anonymous):

yup go on..

OpenStudy (anonymous):

5 - 4x < -7 a. Solve the inequality showing all steps and b. Describe the graph. (Guide: open or closed circle on what number, arrow to the right or left.)

OpenStudy (anonymous):

-4x<-12 x>3 open 0 on 3 and going on towards +ve infinity

OpenStudy (anonymous):

thank you ! (: